OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • 2square root of 6 rational? Yes No

OpenStudy (anonymous):

@Mokeira , @Loser66

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

@Firejay5 , @chlobohoe , @Directrix

OpenStudy (danjs):

Should be \[2\sqrt{3}*2\sqrt{6} = [2*2]*\sqrt{3}\sqrt{6} = 4*\sqrt{3*6} = 4\sqrt{18}\]

OpenStudy (anonymous):

So it is not rational?

OpenStudy (danjs):

root 18 can simplify a bit more, but you cant get rid of the root without making a decimal

OpenStudy (danjs):

\[4\sqrt{18}=4\sqrt{9*2}=4\sqrt{9}\sqrt{2}=4*3\sqrt{2}=12\sqrt{2}\]

OpenStudy (anonymous):

Oh so it is rational right, since it can be simplified?

OpenStudy (danjs):

The most you can simplify it , is to \[12\sqrt{2}\] Is that a rational number?

OpenStudy (anonymous):

I'm going to say yes.

OpenStudy (anonymous):

Is that right??

OpenStudy (danjs):

read through that page.. it tells ya some properties of rational/irrational numbers.,

OpenStudy (anonymous):

Oh okay, it says that the square root of 2 is irrational

OpenStudy (danjs):

also here is more, it even has the square root of 2 as an example... http://www.mathsisfun.com/numbers/irrational-finding.html

OpenStudy (anonymous):

Okay so it is irrational right?

OpenStudy (anonymous):

And thanks for the links, it helped!

OpenStudy (danjs):

yep, welcome